OpenStudy (anonymous):

find the area enclosed between the curves y=3x and y=x^2

OpenStudy (anonymous):

First you need to find the bounds of your integral. Solve, \[3x=x^2\] or \[x^2-3x=0\] which has roots at 0 and 3. (You can find these by the quadratic formula or by inspection) Now, the parabola opens up and the line cuts across it so we need to find:\[\int\limits_{0}^{3} (3x-x^2)dx\] take F(x)= \[\frac{3x^2}{2}-\frac{x^3}{3}\] evaluated F(3)-F(0)=(3/2)(9)-9=9(3/2-1)=9/2

OpenStudy (anonymous):

do i need to draw the graph?

OpenStudy (anonymous):

With these problems drawing the graph helps to see the correct order of subtraction. (In more complicated problems you might have functions that go above and under each other. In this case, you would need to split up the problem into separate integrations to find the total area) As long as the grader understands that you know the order of subtraction or you can just subtract however you like and take the absolute value since the problem asks for area specifically.

OpenStudy (anonymous):

so the answer is final? please help if i was to draw a graph how would i go on succeeding in doing so

OpenStudy (anonymous):

@keith27 Do you still need help?

OpenStudy (anonymous):

yes i do chlorophyll

OpenStudy (anonymous):

they want me to graph it out

OpenStudy (anonymous):

Are you allowed to use graphic calculator?

OpenStudy (anonymous):

or you prefer do manually?

OpenStudy (anonymous):

manually and i dont no wer to begin

OpenStudy (anonymous):

Since y = x² open up with its min O ( 0,0), while y = 3x is straight line go up through 0 (0.0) => the line above y = x².

OpenStudy (anonymous):

Do you know how to draw a straight line???

OpenStudy (anonymous):

yeah, wat bout it

OpenStudy (anonymous):

k den is that all though?

OpenStudy (anonymous):

after drawing straight lines will i be done?

OpenStudy (anonymous):

NO, First you draw BOTH graphs y = x², y = 3x then, observe to see which on above to subtract the other Let the different = 0 => 2 limits Take finite Integral => Area ! DONE

OpenStudy (anonymous):

oh ok il try doing that

OpenStudy (anonymous):

@loggingin has already worked out in details. I just explained the steps!

OpenStudy (anonymous):

ok den thanks